# DST-Adapter: Disentangled-and-Deformable Spatio-Temporal Adapter for Few-shot Action Recognition

### v1

@@ 3.3 End-to-End Adapter Tuning
The learned features by our adapter tuning framework are further used for few-shot action recognition based on the metric-based strategy. It first calculates the frame-wise distance matrix between the input query video and each class prototype derived by simply averaging the corresponding support samples. Then the distance matrix is used to calculate the matching similarity between the query and each class for prediction. The whole model can be optimized in an end-to-end manner using the cross-entropy loss.
@@ 4.1 Experimental Setup
Method | Reference | Backbone | Adaptation | SSv2-Full | SSv2-Small | ||||||
paradigm | 1-shot | 2-shot | 3-shot | 4-shot | 5-shot | 1-shot | 3-shot | 5-shot | |||
CMN [57] | ECCV(18) | ResNet-50 | Full fine-tuning | 34.4 | 43.8 | 33.4 | 42.5 | 46.5 | |||
CMN-J [58] | TPAMI(20) | ResNet-50 | Full fine-tuning | 36.2 | 44.6 | 48.8 | |||||
OTAM [2] | CVPR(20) | ResNet-50 | Full fine-tuning | 42.8 | 49.1 | 51.5 | 52.0 | 52.3 | 36.4 | 45.9 | 48.0 |
ITANet [54] | IJCAI(21) | ResNet-50 | Full fine-tuning | 49.2 | 55.5 | 59.1 | 61.0 | 62.3 | 39.8 | 49.4 | 53.7 |
TRX [26] | CVPR(21) | ResNet-50 | Full fine-tuning | 42.0 | 53.1 | 57.6 | 61.1 | 64.6 | 36.0 | 51.9 | 59.1 |
TAN [22] | AAAI(22) | ResNet-50 | Full fine-tuning | 47.6 | 61.0 | ||||||
MTFAN [44] | CVPR(22) | ResNet-50 | Full fine-tuning | 45.7 | 60.4 | ||||||
STRM [34] | CVPR(22) | ResNet-50 | Full fine-tuning | 43.1 | 53.3 | 59.1 | 61.7 | 68.1 | 37.1 | 49.2 | 55.3 |
HyRSM [40] | CVPR(22) | ResNet-50 | Full fine-tuning | 54.3 | 62.2 | 65.1 | 67.9 | 69.0 | 40.6 | 52.3 | 56.1 |
Nguyen et al. [24] | ECCV(22) | ResNet-50 | Full fine-tuning | 43.8 | 61.1 | ||||||
Huang et al. [15] | ECCV(22) | ResNet-50 | Full fine-tuning | 49.3 | 66.7 | 38.9 | 61.6 | ||||
HCL [55] | ECCV(22) | ResNet-50 | Full fine-tuning | 47.3 | 54.5 | 59.0 | 62.4 | 64.9 | 38.7 | 49.1 | 55.4 |
SloshNet [47] | AAAI(23) | ResNet-50 | Full fine-tuning | 46.5 | 68.3 | ||||||
GgHM [48] | ICCV(23) | ResNet-50 | Full fine-tuning | 54.5 | 69.2 | ||||||
MoLo [42] | CVPR(23) | ResNet-50 | Full fine-tuning | 56.6 | 62.3 | 67.0 | 68.5 | 70.6 | 42.7 | 52.9 | 56.4 |
ST-Adapter [25] | NeurIPS(22) | ResNet-50 | Adapter tuning | 52.2 | 59.9 | 64.1 | 67.1 | 68.7 | 41.9 | 51.5 | 55.7 |
DST-Adapter (Ours) | ResNet-50 | Adapter tuning | 57.0 | 65.2 | 69.5 | 71.4 | 73.6 | 45.8 | 56.6 | 60.9 | |
CLIP-FSAR [41] | ArXiv(23) | CLIP-ViT-B | Full fine-tuning | 61.9 | 64.9 | 68.1 | 70.9 | 72.1 | 54.5 | 58.6 | 61.8 |
ST-Adapter [25] | NeurIPS(22) | CLIP-ViT-B | Adapter tuning | 64.2 | 72.4 | 76.0 | 77.4 | 79.5 | 53.1 | 64.2 | 68.0 |
DST-Adapter (Ours) | CLIP-ViT-B | Adapter tuning | 66.7 | 75.3 | 78.3 | 80.1 | 81.9 | 55.0 | 65.8 | 69.3 |
@@
Pre-trained backbone. Most prior arts employs the ResNet-50 [13] pre-trained on ImageNet [5] as backbone to extract frame features. Recently, CLIP-FSAR [41] conduct experiments on CLIP-ViT-B [7] and get remarkable performance. To verify the effectiveness of our DST-Adapter, we instantiate our method with both pre-trained backbones.
@@
Implementation details. We insert one DST-Adapter module in each stage of the pre-trained model (4 stages for ResNet-50 and 12 stages for CLIP-ViT-B) as shown in Figure 2 (a). We tune the sampling kernel for DSTA-S and DSTA-T for each backbone and the details are provided in the supplementary material. For a fair comparison with previous methods [2, 40, 42], we uniformly sample 8 frames (i.e., ) as the input of a video. We employ several standard techniques for data augmentation such as random crop and color jitter in the training stage. Bi-MHM [40] is used as the matching metric in comparison with other methods. Besides, OTAM [2] and TRX [26] are additionally used in the ablation study. We train the model episodically with Adam [18] optimizer. Following the common practice [40, 42], in the test stage, we randomly select 10,000 episodes from the test set and report the average accuracy.
@@ 4.2 Comparison with State-of-the-Art Methods
Benchmarks sensitive to temporal features. Table 1 shows the comparative results of different methods on SSv2-Full and SSv2-Small, which are quite challenging since both the temporal and spatial features are crucial to the action recognition. Our DST-Adapter achieves the best performance and outperforms other methods by a large margin with both backbones in all few-shot settings except ‘5-shot’ on SSv2-Small. These results demonstrate the advantages of our method over other methods in dealing with challenging scenarios where the temporal features are critical for action recognition. In particular, our DST-Adapter outperforms ST-Adapter substantially on all settings, which reveals the benefit of two core technical designs of DST-Adapter, namely the dual-pathway architecture for disentangled learning of spatial and temporal features and Deformable Spatio-Temporal Attention for learning features in a global view.
Method | Reference | Backbone | Adaptation | Kinetics | HMDB51 | UCF101 | ||||||||
paradigm | 1-shot | 2-shot | 3-shot | 4-shot | 5-shot | 1-shot | 3-shot | 5-shot | 1-shot | 3-shot | 5-shot | |||
CMN [57] | ECCV(18) | ResNet-50 | Full fine-tuning | 57.3 | 67.5 | 72.5 | 74.7 | 76.0 | ||||||
CMN-J [58] | TPAMI(20) | ResNet-50 | Full fine-tuning | 60.5 | 70.0 | 75.6 | 77.3 | 78.9 | ||||||
OTAM [2] | CVPR(20) | ResNet-50 | Full fine-tuning | 73.0 | 75.9 | 78.7 | 81.9 | 85.8 | 54.5 | 65.7 | 68.0 | 79.9 | 87.0 | 88.9 |
ITANet [54] | IJCAI(21) | ResNet-50 | Full fine-tuning | 73.6 | 84.3 | |||||||||
TRX [26] | CVPR(21) | ResNet-50 | Full fine-tuning | 63.6 | 76.2 | 81.8 | 83.4 | 85.9 | 53.1 | 66.8 | 75.6 | 78.2 | 92.4 | 96.1 |
TAN [22] | AAAI(22) | ResNet-50 | Full fine-tuning | 72.8 | 85.8 | 59.7 | 73.9 | 81.9 | 95.1 | |||||
MTFAN [44] | CVPR(22) | ResNet-50 | Full fine-tuning | 74.6 | 87.4 | 59.0 | 74.6 | 84.8 | 95.1 | |||||
STRM [34] | CVPR(22) | ResNet-50 | Full fine-tuning | 62.9 | 76.4 | 81.1 | 83.8 | 86.7 | 52.3 | 67.4 | 77.3 | 80.5 | 92.7 | 96.9 |
HyRSM [40] | CVPR(22) | ResNet-50 | Full fine-tuning | 73.7 | 80.0 | 83.5 | 84.6 | 86.1 | 60.3 | 71.7 | 76.0 | 83.9 | 93.0 | 94.7 |
Nguyen et al. [24] | ECCV(22) | ResNet-50 | Full fine-tuning | 74.3 | 87.4 | 59.6 | 76.9 | 84.9 | 95.9 | |||||
Huang et al. [15] | ECCV(22) | ResNet-50 | Full fine-tuning | 73.3 | 86.4 | 60.1 | 77.0 | 71.4 | 91.0 | |||||
HCL [55] | ECCV(22) | ResNet-50 | Full fine-tuning | 73.7 | 79.1 | 82.4 | 84.0 | 85.8 | 59.1 | 71.2 | 76.3 | 82.5 | 91.0 | 93.9 |
SloshNet [47] | AAAI(23) | ResNet-50 | Full fine-tuning | 70.4 | 87.0 | 59.4 | 77.5 | 86.0 | 97.1 | |||||
GgHM [48] | ICCV(23) | ResNet-50 | Full fine-tuning | 74.9 | 87.4 | 61.2 | 76.9 | 85.2 | 96.3 | |||||
MoLo [42] | CVPR(23) | ResNet-50 | Full fine-tuning | 74.0 | 80.4 | 83.7 | 84.7 | 85.6 | 60.8 | 72.0 | 77.4 | 86.0 | 93.5 | 95.5 |
ST-Adapter [25] | NeurIPS(22) | ResNet-50 | Adapter tuning | 73.0 | 79.9 | 82.8 | 84.6 | 85.1 | 60.3 | 71.4 | 74.7 | 84.6 | 92.9 | 94.5 |
DST-Adapter (Ours) | ResNet-50 | Adapter tuning | 75.0 | 81.4 | 84.3 | 85.8 | 87.0 | 61.6 | 73.0 | 76.6 | 86.9 | 94.4 | 95.6 | |
CLIP-FSAR [41] | ArXiv(23) | CLIP-ViT-B | Full fine-tuning | 89.7 | 92.9 | 94.2 | 94.8 | 95.0 | 75.8 | 84.1 | 87.7 | 96.6 | 98.4 | 99.0 |
ST-Adapter [25] | NeurIPS(22) | CLIP-ViT-B | Adapter tuning | 88.5 | 92.6 | 94.0 | 94.6 | 95.1 | 74.1 | 84.3 | 87.3 | 95.9 | 98.5 | 98.9 |
DST-Adapter (Ours) | CLIP-ViT-B | Adapter tuning | 89.3 | 93.1 | 94.4 | 95.2 | 95.5 | 77.1 | 86.4 | 88.2 | 96.4 | 98.7 | 99.1 |
@@
To obtain more insight into the strength and weakness of our DST-Adapter, we select four categories of actions that our method excels at as well as four categories that our model struggles with at compared with the state-of-the-art MoLo and ST-Adapter. Figure 3 shows that our model achieves large performance superiority over other methods on the relatively complex actions requiring careful reasoning via learning temporal features for recognition, whilst performing poorly on simple actions that can recognized primarily based on spatial features. Such results are consistent with the previous results that our model excels at recognizing actions involving complex temporal dynamics.
@@
Benchmarks relying on spatial features. Table 2 presents the experimental results on the benchmarks relying more on spatial features, including Kinetics, HMDB51 and UCF101. Our method still performs best on most of the settings, which manifest the effectiveness and robustness of our method in these scenarios. We also observe that the performance improvement by our DST-Adapter over other method, especially ST-Adapter, is smaller than that on SSv2 datasets. It is reasonable since the action recognition in these datasets relies less on the temporal features and thus the disentangled feature encoding of our method yields limited performance gain.
@@
Comparison of efficiency. We compare the efficiency between our DST-Adapter and other methods in terms of tunable parameter size and memory usage in Figure 4. It shows that the tunable parameter size of both our model and ST-Adapter is significantly smaller than other methods that are based on full fine-tuning for task adaptation, which reveals the advantage of the adapter tuning paradigm over the full fine-tuning paradigm for task adaptation. Meanwhile, our method outperforms these full fine-tuning based methods substantially while enjoying much fewer tunable parameters, which validates the superiority of our method.
@@ 4.3 Ablation study
We conduct ablation study by using ResNet-50 as the pre-trained backbone and Bi-MHM [40] as the matching metric.
Comparison between different Adapters. We first compare the performance of different Adapters in Table 3. Vanilla-Adapter is the initially proposed adapter [14] which only contains the downsampling and upsampling layers with a GELU nonlinearity in between. As described in Section 3.2, DST-Adapter is the convolutional version of our DST-Adapter which follows the dual-pathway structure but models spatial and temporal pathways with 3D convolution. The performance of full fine-tuning is also provided for reference. We make following observations.
@@
Effect of the proposed Deformable Spatial-Temporal Attention (DSTA). Comparing our DST-Adapter with DST-Adapter, the proposed DSTA yields a large performance gain, especially on the SSv2-Full dataset, which is more challenging due to the necessity of encoding temporal features.
@@
DSTA-Uniform | 55.8 | 72.2 | 73.9 | 86.0 |
DSTA-S & DSTA-T | 57.0 | 73.6 | 75.0 | 87.0 |
Effect of configuration of the sampling kernel of DSTA. To investigate the effect of configuring specialized sampling kernels for DSTA-S in the spatial pathway and DSTA-T in the temporal pathway respectively, we additionally configure a DSTA-Uniform module, which performs sampling with uniform density in both the spatial and temporal domains. Then we compare the performance of our DST-Adapter using two different settings, including 1) ‘DSTA-Uniform’, in which DSTA-Uniform is used to model both the spatial and temporal pathways and 2) ‘DSTA-S & DSTA-T’, in which DST-Adapter uses DSTA-S and DSTA-T to model the spatial and temporal pathways respectively. As shown in Table 4, the distinct performance gap between these two settings reveals the advantage of configuring specialized sampling kernels for the spatial and temporal pathways.
Generalization across different matching metrics. To evaluate the generalization of our DST-Adapter across different matching metrics, we conduct the experiments using three classical matching metrics on SSv2-Full and Kinetics, including OTAM [2], TRX [26], and Bi-MHM [40], and compare our model with the state-of-the-art methods corresponding to each metric. Table 5 presents the results, which show that our DST-Adapter consistently outperforms other methods, especially on challenging SSv2-Full benchmark, validating the well generalization of the proposed DST-Adapter across different matching metrics.
@@
DST-Adapter | 57.0 | 73.6 | 75.0 | 87.0 |
4.4 Visualization
In this section we visualize the shifted reference points learned by our DST-Adapter to investigate the effectiveness of the proposed DSTA qualitatively. Following Deformable Attention [46], for each shifted reference point serving as key and value, we accumulate the attention weights assigned by all queries as its relative importance. Then we visualize top-100 most important shifted reference points in each pathway for a video sample, as shown in Figure 5. Note that we sample 8 frames for all videos, if a reference point is shifted to a temporal position between two sampled frames, we visualize it on the nearest frame approximately for simplicity.
The results show that the shifted reference points learned by our model can always focus on the salient targets that are critical for action recognition in both the spatial and temporal pathways. Another interesting observation is that the shifted points in the spatial pathway are distributed densely in some specific frames to capture the spatial appearance features for the targets, while the shifted points in the temporal pathway are distributed uniformly among all frames to capture the temporal dynamic features. This is consistent with the design of specialized tailored DSTA modules for different pathways by configuring the sampling kernels accordingly.
5 Conclusion
We present DST-Adapter, which is a novel adapter tuning method for few-shot action recognition. It is designed in dual-pathway architecture, which allows for disentangled encoding of spatial and temporal features. Moreover, we specially design the Deformable Spatio-Temporal Attention (DSTA) as the essential component of DST-Adapter, which enables it to encode features in a global view while maintaining lightweight design. Extensive experiments demonstrate that our model compares favorably with other state-of-the-art methods, especially in the challenging scenarios involving complex temporal features.
Supplementary Material
Appendix A Instantiation on ResNet-18 and ResNet-34
To investigate the performance of our DST-Adapter on smaller backbones, we instantiate our model with ResNet-18 and ResNet-34 pre-trained on ImageNet, and conduct experiments to compare our model with other methods under the same experimental settings. The results in Table 8 show that our DST-Adapter achieves the best performance using both backbones in all settings, which demonstrates the robustness of our method. While DST-Adapter, the convolutional version of our model, also performs well, it is still inferior to DST-Adapter, which manifests the effectiveness of the proposed deformable spatio-temporal attention (DSTA).
Appendix B Details of the Tuned Sampling Kernels
The sampling kernels of DSTA-S in the spatial pathway and DSTA-T in the temporal pathway of our DST-Adapter can be tuned on a held-out validation set. Generally, DSTA-S should sample denser reference points in the spatial domain while DSTA-T samples more points in the temporal domain. For the instantiation of our model with ResNet, we configure the sampling kernel of DSTA in each convolutional stage individually since the feature maps in different convolutional stages may have different size. In contrast, we only need to tune one configuration for sampling kernel when using ViT as the backbone since the feature map always has fixed size in different stages. Table 6 shows the tuned configurations of the sampling kernels of DSTA-S and DSTA-T with different backbones. Besides, we also provide the configurations of sampling kernel for DSTA-Uniform module which is constructed to validate the effect of configuring the sampling kernel of DSTA in Table 4.
in DSTA-S | in DSTA-T | in DSTA-Uniform | ||
ResNet-50 | (8, 56, 56) | (2, 8, 8) | (8, 4, 4) | (4, 4, 4) |
ResNet-50 | (8, 28, 28) | (2, 4, 4) | (8, 2, 2) | (4, 4, 4) |
ResNet-50 | (8, 14, 14) | (2, 4, 4) | (8, 2, 2) | (4, 4, 4) |
ResNet-50 | (8, 7, 7) | (2, 2, 2) | (8, 1, 1) | (4, 4, 4) |
CLIP-ViT-B | (8, 14, 14) | (2, 4, 4) | (8, 2, 2) | (4, 4, 4) |
Appendix C Effect of the Inserted Position of DST-Adapter
Theoretically, our DST-Adapter can be inserted into any position of the backbone flexibly. To investigate the effect of the inserted position of DST-Adapter on the performance of the model, we conduct experiments with four different ways of inserting DST-Adapters into the pre-trained CLIP-ViT-B backbone (comprising 12 learning stages) on SSv2-Small dataset: a) early-insertion, which inserts the DST-Adapter into each of first 6 stages (close to the input), b) late-insertion that inserts the DST-Adapter into each of last 6 stages (close to the output), c) skip-insertion, which inserts the adapter into the backbone every two stages and d) full-insertion that inserts the adapter into each learning stage.
As shown in Table 7, late-insertion, namely inserting the proposed DST-Adapters into the last 6 stages, yields better performance than early-insertion, which implies that adapter tuning is more effective for task adaptation in deeper layers than in the shallower layers. It is reasonable since deeper layers generally capture the high-level semantic features, which are more relevant to task adaptation. Besides, full-insertion achieves the best performance at the expense of slightly more tunable parameters and memory usage.
Early-insertion | 4.1% | 15.5GB | 48.4 | 64.6 |
Late-insertion | 4.1% | 15.2GB | 54.2 | 68.5 |
Skip-insertion | 4.1% | 15.2GB | 53.3 | 67.9 |
Full-insertion | 7.9% | 16.7GB | 55.0 | 69.3 |
Appendix D Effect of the Bottleneck Ratio of DST-Adapter
The tunable parameter size is mainly determined by the bottleneck ratio of DST-Adapter, defined as the ratio of downsampled channel numbers to the initial size. Thus, we can balance between the model efficiency in terms of tunable parameter size and model effectiveness in terms of classification accuracy by tuning the bottleneck ratio. As shown in Table 9, larger bottleneck ratios typically yield more performance improvement while introducing more tunable parameters, and we set the bottleneck ratio to 0.25 in all the experiments based on the results.

### v2

@@ 3.3 End-to-End Adapter Tuning
The learned features by our adapter tuning framework are further used for few-shot action recognition based on the metric-based strategy. It first calculates the frame-wise distance matrix between the input query video and each class prototype derived by simply averaging the corresponding support samples. Then the distance matrix is used to calculate the matching similarity between the query and each class for prediction. Formally, the similarity between query and class prototype can be expressed as:
where and denote the features of the -th frame in query and the -th frame in class prototype while is the matching metric such as OTAM [2] and Bi-MHM [44]. The whole model can be optimized in an end-to-end manner using the cross-entropy loss.
@@ 4.1 Experimental Setup
Pre-trained backbone. Most prior arts employs the ResNet-50 [13] pre-trained on ImageNet [5] as backbone to extract frame features. Recently, CLIP-FSAR [42] conduct experiments on CLIP-ViT-B [7] and get remarkable performance. To verify the effectiveness of our D2ST-Adapter, we instantiate our method with both pre-trained backbones.
@@
Implementation details. We insert one D2ST-Adapter module in each stage of the pre-trained model (4 stages for ResNet-50 and 12 stages for CLIP-ViT-B) as shown in Figure 2 (a). We tune the sampling kernel for aDSTA-S and aDSTA-T for each backbone and the details are provided in the supplementary material. For a fair comparison with previous methods [2, 44, 43], we uniformly sample 8 frames (i.e., ) as the input of a video. We employ several standard techniques for data augmentation such as random crop and color jitter in the training stage. Bi-MHM [44] is used as the matching metric in comparison with other methods. Besides, OTAM [2] and TRX [28] are additionally used in the ablation study. We train the model episodically with Adam [18] optimizer. Following the common practice [44, 43], in the test stage, we randomly select 10,000 episodes from the test set and report the average accuracy.
Method | Backbone | Adaptation | SSv2-Full | SSv2-Small | ||||||
paradigm | 1-shot | 2-shot | 3-shot | 4-shot | 5-shot | 1-shot | 3-shot | 5-shot | ||
CMN [59] | ResNet-50 | Full fine-tuning | 34.4 | 43.8 | 33.4 | 42.5 | 46.5 | |||
CMN-J [60] | ResNet-50 | Full fine-tuning | 36.2 | 44.6 | 48.8 | |||||
OTAM [2] | ResNet-50 | Full fine-tuning | 42.8 | 49.1 | 51.5 | 52.0 | 52.3 | 36.4 | 45.9 | 48.0 |
ITANet [56] | ResNet-50 | Full fine-tuning | 49.2 | 55.5 | 59.1 | 61.0 | 62.3 | 39.8 | 49.4 | 53.7 |
TRX [28] | ResNet-50 | Full fine-tuning | 42.0 | 53.1 | 57.6 | 61.1 | 64.6 | 36.0 | 51.9 | 59.1 |
TA2N [22] | ResNet-50 | Full fine-tuning | 47.6 | 61.0 | ||||||
MTFAN [46] | ResNet-50 | Full fine-tuning | 45.7 | 60.4 | ||||||
STRM [36] | ResNet-50 | Full fine-tuning | 43.1 | 53.3 | 59.1 | 61.7 | 68.1 | 37.1 | 49.2 | 55.3 |
HyRSM [44] | ResNet-50 | Full fine-tuning | 54.3 | 62.2 | 65.1 | 67.9 | 69.0 | 40.6 | 52.3 | 56.1 |
Nguyen et al. [24] | ResNet-50 | Full fine-tuning | 43.8 | 61.1 | ||||||
Huang et al. [15] | ResNet-50 | Full fine-tuning | 49.3 | 66.7 | 38.9 | 61.6 | ||||
HCL [57] | ResNet-50 | Full fine-tuning | 47.3 | 54.5 | 59.0 | 62.4 | 64.9 | 38.7 | 49.1 | 55.4 |
SloshNet [49] | ResNet-50 | Full fine-tuning | 46.5 | 68.3 | ||||||
GgHM [50] | ResNet-50 | Full fine-tuning | 54.5 | 69.2 | ||||||
MoLo [43] | ResNet-50 | Full fine-tuning | 56.6 | 62.3 | 67.0 | 68.5 | 70.6 | 42.7 | 52.9 | 56.4 |
ST-Adapter [25] | ResNet-50 | Adapter tuning | 52.2 | 59.9 | 64.1 | 67.1 | 68.7 | 41.9 | 51.5 | 55.7 |
D2ST-Adapter (Ours) | ResNet-50 | Adapter tuning | 57.0 | 65.2 | 69.5 | 71.4 | 73.6 | 45.8 | 56.6 | 60.9 |
CLIP-FSAR [42] | CLIP-ViT-B | Full fine-tuning | 61.9 | 64.9 | 68.1 | 70.9 | 72.1 | 54.5 | 58.6 | 61.8 |
AIM [51] | CLIP-ViT-B | Adapter tuning | 63.7 | 72.0 | 75.2 | 77.3 | 79.2 | 52.8 | 64.0 | 67.5 |
DUALPATH [26] | CLIP-ViT-B | Adapter tuning | 64.5 | 72.6 | 76.0 | 78.0 | 79.8 | 53.5 | 64.4 | 68.1 |
ST-Adapter [25] | CLIP-ViT-B | Adapter tuning | 64.2 | 72.4 | 76.0 | 77.4 | 79.5 | 53.1 | 64.2 | 68.0 |
D2ST-Adapter (Ours) | CLIP-ViT-B | Adapter tuning | 66.7 | 75.3 | 78.3 | 80.1 | 81.9 | 55.0 | 65.8 | 69.3 |
@@ 4.2 Comparison with State-of-the-Art Methods
Benchmarks sensitive to temporal features. Table 1 shows the comparative results of different methods on SSv2-Full and SSv2-Small, which are quite challenging since both the temporal and spatial features are crucial to the action recognition. Our D2ST-Adapter achieves the best performance and outperforms other methods by a large margin with both backbones in all few-shot settings except ‘5-shot’ on SSv2-Small. These results demonstrate the advantages of our method over other methods in dealing with challenging scenarios where the temporal features are critical for action recognition. In particular, our D2ST-Adapter outperforms all existing adapter tuning methods for video data, including AIM, DUALPATH, and ST-Adapter, substantially on all settings, which reveals the benefit of two core technical designs of D2ST-Adapter, namely the dual-pathway architecture for disentangled learning of spatial and temporal features and anisotropic Deformable Spatio-Temporal Attention (aDSTA) for learning features in a global view.
@@
To gain more insight into the strengths and weaknesses of our D2ST-Adapter, we select four categories of actions where our method excels, as well as four categories where our model struggles in comparison to the state-of-the-art MoLo and ST-Adapter. Figure 3 shows that our model achieves large performance superiority over other methods on the relatively complex actions requiring careful reasoning via learning temporal features for recognition, whilst performing poorly on simple actions that can be recognized primarily based on spatial features. Such results are consistent with the previous results that our model excels at recognizing actions involving complex temporal dynamics.
Method | Backbone | Adaptation | Kinetics | HMDB51 | UCF101 | ||||||||
paradigm | 1-shot | 2-shot | 3-shot | 4-shot | 5-shot | 1-shot | 3-shot | 5-shot | 1-shot | 3-shot | 5-shot | ||
CMN [59] | ResNet-50 | Full fine-tuning | 57.3 | 67.5 | 72.5 | 74.7 | 76.0 | ||||||
CMN-J [60] | ResNet-50 | Full fine-tuning | 60.5 | 70.0 | 75.6 | 77.3 | 78.9 | ||||||
OTAM [2] | ResNet-50 | Full fine-tuning | 73.0 | 75.9 | 78.7 | 81.9 | 85.8 | 54.5 | 65.7 | 68.0 | 79.9 | 87.0 | 88.9 |
ITANet [56] | ResNet-50 | Full fine-tuning | 73.6 | 84.3 | |||||||||
TRX [28] | ResNet-50 | Full fine-tuning | 63.6 | 76.2 | 81.8 | 83.4 | 85.9 | 53.1 | 66.8 | 75.6 | 78.2 | 92.4 | 96.1 |
TA2N [22] | ResNet-50 | Full fine-tuning | 72.8 | 85.8 | 59.7 | 73.9 | 81.9 | 95.1 | |||||
MTFAN [46] | ResNet-50 | Full fine-tuning | 74.6 | 87.4 | 59.0 | 74.6 | 84.8 | 95.1 | |||||
STRM [36] | ResNet-50 | Full fine-tuning | 62.9 | 76.4 | 81.1 | 83.8 | 86.7 | 52.3 | 67.4 | 77.3 | 80.5 | 92.7 | 96.9 |
HyRSM [44] | ResNet-50 | Full fine-tuning | 73.7 | 80.0 | 83.5 | 84.6 | 86.1 | 60.3 | 71.7 | 76.0 | 83.9 | 93.0 | 94.7 |
Nguyen et al. [24] | ResNet-50 | Full fine-tuning | 74.3 | 87.4 | 59.6 | 76.9 | 84.9 | 95.9 | |||||
Huang et al. [15] | ResNet-50 | Full fine-tuning | 73.3 | 86.4 | 60.1 | 77.0 | 71.4 | 91.0 | |||||
HCL [57] | ResNet-50 | Full fine-tuning | 73.7 | 79.1 | 82.4 | 84.0 | 85.8 | 59.1 | 71.2 | 76.3 | 82.5 | 91.0 | 93.9 |
SloshNet [49] | ResNet-50 | Full fine-tuning | 70.4 | 87.0 | 59.4 | 77.5 | 86.0 | 97.1 | |||||
GgHM [50] | ResNet-50 | Full fine-tuning | 74.9 | 87.4 | 61.2 | 76.9 | 85.2 | 96.3 | |||||
MoLo [43] | ResNet-50 | Full fine-tuning | 74.0 | 80.4 | 83.7 | 84.7 | 85.6 | 60.8 | 72.0 | 77.4 | 86.0 | 93.5 | 95.5 |
ST-Adapter [25] | ResNet-50 | Adapter tuning | 73.0 | 79.9 | 82.8 | 84.6 | 85.1 | 60.3 | 71.4 | 74.7 | 84.6 | 92.9 | 94.5 |
D2ST-Adapter (Ours) | ResNet-50 | Adapter tuning | 75.0 | 81.4 | 84.3 | 85.8 | 87.0 | 61.6 | 73.0 | 76.6 | 86.9 | 94.4 | 95.6 |
CLIP-FSAR [42] | CLIP-ViT-B | Full fine-tuning | 89.7 | 92.9 | 94.2 | 94.8 | 95.0 | 75.8 | 84.1 | 87.7 | 96.6 | 98.4 | 99.0 |
AIM [51] | CLIP-ViT-B | Adapter tuning | 88.4 | 92.7 | 94.2 | 94.6 | 95.3 | 74.2 | 83.8 | 86.9 | 95.4 | 98.2 | 98.5 |
DUALPATH [26] | CLIP-ViT-B | Adapter tuning | 88.8 | 92.8 | 94.3 | 94.7 | 95.4 | 74.9 | 84.6 | 87.5 | 95.7 | 98.4 | 98.7 |
ST-Adapter [25] | CLIP-ViT-B | Adapter tuning | 88.5 | 92.6 | 94.0 | 94.6 | 95.1 | 74.1 | 84.3 | 87.3 | 95.9 | 98.5 | 98.9 |
D2ST-Adapter (Ours) | CLIP-ViT-B | Adapter tuning | 89.3 | 93.1 | 94.4 | 95.2 | 95.5 | 77.1 | 86.4 | 88.2 | 96.4 | 98.7 | 99.1 |
@@
Benchmarks relying on spatial features. Table 2 presents the experimental results on the benchmarks relying more on spatial features, including Kinetics, HMDB51 and UCF101. Our method still performs best on most of the settings, which manifest the effectiveness and robustness of our method in these scenarios. We also observe that the performance improvement by our D2ST-Adapter over other method, especially ST-Adapter, is smaller than that on SSv2 datasets. It is reasonable since the action recognition in these datasets relies less on the temporal features and thus the disentangled feature encoding of our method yields limited performance gain.
@@
Comparison of efficiency. We compare the efficiency between our D2ST-Adapter and other methods in terms of tunable parameter size and memory usage in Figure 4. It shows that the tunable parameter size of both our model and ST-Adapter is significantly smaller than other methods that are based on full fine-tuning, which reveals the advantage of the adapter tuning paradigm over the full fine-tuning paradigm for task adaptation. Meanwhile, our method outperforms these full fine-tuning based methods substantially while enjoying much fewer tunable parameters, which validates the superiority of our method.
@@ 4.3 Ablation study
We conduct ablation study on SSv2-Full and Kinetics datasets, using ResNet-50 as the pre-trained backbone and Bi-MHM [44] as the matching metric.
Comparison between different Adapters. We first compare the performance of different Adapters in Table 3. As shown in Figure 1, Vanilla-Adapter [14] only contains the downsampling and upsampling layers with a GELU nonlinearity in between. As described in Section 3.2, DST-Adapter is the convolutional version of our D2ST-Adapter which follows the dual-pathway structure but models spatial and temporal pathways with 3D convolution. The performance of full fine-tuning is also provided for reference. We make following observations.
@@
Effect of the proposed anisotropic Deformable Spatial-Temporal Attention (aDSTA). Comparing our D2ST-Adapter with DST-Adapter, the proposed aDSTA yields a large performance gain, especially on the SSv2-Full dataset, which is more challenging due to the necessity of encoding temporal features.
@@
aDSTA-Uniform | 55.8 | 72.2 | 73.9 | 86.0 |
aDSTA-S & aDSTA-T | 57.0 | 73.6 | 75.0 | 87.0 |
@@
D2ST-Adapter | 57.0 | 73.6 | 75.0 | 87.0 |
Effect of configuration of the sampling kernel of aDSTA. To investigate the effect of configuring specialized sampling kernels for aDSTA-S in the spatial pathway and aDSTA-T in the temporal pathway respectively, we additionally configure an aDSTA-Uniform module, which performs sampling with uniform density in both the spatial and temporal domains. Then we compare the performance of our D2ST-Adapter using two different settings, including 1) ‘aDSTA-Uniform’, in which aDSTA-Uniform is used to model both the spatial and temporal pathways and 2) ‘aDSTA-S & aDSTA-T’, in which D2ST-Adapter uses aDSTA-S and aDSTA-T to model the spatial and temporal pathways respectively. As shown in Table 4, the distinct performance gap between these two settings reveals the advantage of configuring specialized sampling kernels for the spatial and temporal pathways.